) Write a pseudocode that calculates and prints the average of
several integers. Assume the last value read with scanf is the sentinel 9999. A typical input sequence
might be
10 8 11 7 9 9999
indicating that the average of all the values preceding 9999 is to be calculated.
1
Expert's answer
2015-12-22T04:44:17-0500
Solution
Declare a as integer
Declare sum as real
Declare avg as real
Declare n as integer
Let n=0
Let sum=0
Print ‘Input set of numbers, 9999 for end of sequence ’

repeat
Input a
if a≠9999 then
Let n=n+1
Les sum=sum+a
end if
until a≠9999

if n≠0 then
avg=sum/n
Print ‘Average is ’, avg
Else
Print ‘Sequence is empty’
End if